OpenStudy (anonymous):

If h(x) = −2x − 10, find h(−4). −2 −18 −3 −16

OpenStudy (diyadiya):

Just plug in -4 in place of x :)

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

-2?

OpenStudy (diyadiya):

Nope!

OpenStudy (diyadiya):

-2(-4)-11=8-11= ?

OpenStudy (anonymous):

-3?

OpenStudy (diyadiya):

Right!